'R' and 's' are partners RS. 40000 and 3000 respectively. And sharing in the ratio of 3:2. This ratio equation please????





Answer by satyareddy22(84)

R-40000----3=>(40000)3=120000
S-3000-----2=>(3000)2=6000
R/S = 120000/6000= 20/1
R:S =20:1